On a journey from planet P to Venus a rocket is traveling at a constant speed. During the journey the rocket Travels past the moon in 4.8x10^5 km.
i) Find the distance travelled by the Rocket in 12 days and give your answer in standard form.
ii) Given that the distance between planet P and Venus is 4.8x10^7 km, find the time taken in days for the journey.

Found 2 solutions by Alex.33, ikleyn:
Answer by Alex.33(110)

#warning: "During the journey the rocket Travels past the moon in 4.8x10^5 km."
This description is grammatically incorrect and it's impossible to determine what you had to say. My assumption is that perhaps you mean "4.8*10^5" is the speed, (maybe in km/h and you left "/h" behind), or you're just describing another concept.

if that means "During the journey the rocket Travels past the moon in 4.8*10^5 km/h." I'll go ahead and do the following:
i)Ans=12*24*4.8*10^5=138240000=1.3824*10^8 km.
ii)Ans=4.8*10^7/4.8*10^5=100 days.

if you mean by it other stuffs, sorry for not being able to help. You can make your words clearer and submit the question again, and I'm pretty sure you'll get a good response.

Hope that helps!

Answer by ikleyn(52781)

.
How this problem is formulated, worded, presented and posted, it MAKES NO SENSE and can not be solved.

Making guesses regarding the condition is not the method solving Math problems.

--------------

By the way, teaching students to express their thoughts and statements in correct form and in correct terms
is essential part of the mathematical education.